When a 70 kg person climbs into a 1200 kg car, the car’s springs compress vertically by 0.025 m. What will be the frequency of vibration when the car hits a bump? (Hint: Find k first.)      Ignore damping.              
Answer #1

F = mxg = 70x9.8= 686 N 686 kx = 686 N = K² 686 = 27440 N/m 0.005 Total mass = 1200+70 = 1270 kg. Angular frequency, we 4.65